Komiko Tanaka invests $19,000 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 10 percent before-tax rate of return. She plans to invest for the long term. a. How much cash will Komiko retain, after-taxes, if she holds the investment for 5 years and then she sells it when the long-term capital gains rate is 15 percent? (Do not round your intermediate calculations. Round your final answer to the nearest whole dollar amount) $ 28,860 b. What is Komikos after-tax rate of return on her investment in part (a)? (Do not round your intermediate celculations. Round your percentage answer to 2 decimal places.) tax rate 872 percent
gains rate is 25 percent? (Do not round your intermediate colculations. Round your final answer to the nearest whole dollar amount.) geventasgk answer ton her investment i part ic? (o not round your intermediate calculations. Round your percentage answer to 2 decimal places) percent
e. How much cash will Komiko retain, after taxes, if she holds the investment for 15 years and then she sells when the long-term capital gains rate is 15 percent? (Do not round your intermediate calculations. Round your final answer to the nearest whole dollar amount.) t. What is Komikos after-tax rate of return on her investment in part (e? (Do not round your intermediate calculetions. Round your percentage answer to 2 decimal places.) percent After-tax rate of return
Answer #1

a)

Amount received $ 30,600.00
Basis in stock $ 19,000.00
long-term capital gain $ 11,600.00
Capital Gain Rate 15%
Tax on Gain $    1,740.00

Cash retained = $30,600 – 1,740 = $28,860

b) (28,860/19,000)1/5 -1 = 8.7%

c)

Amount received $   30,600.00
Basis in stock $   19,000.00
long-term capital gain $   11,600.00
Capital Gain Rate 25%
Tax on Gain $      2,900.00

Cash retained = $30,600 – 2,900 = $27,700

d) ($27,700/19,000)1/5– 1 = 7.83 %

e)

Amount received $   79,368.00
Basis in stock $   19,000.00
long-term capital gain $   60,368.00
Capital Gain Rate 15%
Tax on Gain $      9,055

Cash retained = $79,368 – 9055 = $70,313

f)($70,313/19,000)1/15– 1 = 9.11 percent